OpenStudy (anonymous):

PreCalc & Trig Help! Simplify the expression using a sum or difference formula: sin(x+y)sec x sec y

OpenStudy (anonymous):

^opps, show all steps too please.

OpenStudy (anonymous):

\[\sin(x+y)=\sin(x)\cos(y)+\cos(x)\sin(y)\] then multiply out

OpenStudy (anonymous):

how do you get to that point? like, can you start from the beginning because I need to show all steps for full credit.

OpenStudy (anonymous):

you can look in your text for a proof of that result, but that is not what you are being asked to do. you are being asked to multiply out, the point of the exercise it to use that identity, not to derive it

OpenStudy (anonymous):

in other words, you are being asked a question to see if you know the identity multiplying out is now a matter of algebra

OpenStudy (anonymous):

I just don't understand how you went from sin(x+y)sec(x)sec(y) to that..

OpenStudy (anonymous):

oh i have confused you that is not the answer at all that is the identity \[\sin(x+y)=\sin(x)\cos(y)+\cos(x)\sin(y)\]

OpenStudy (anonymous):

Yes, you have confused me..

OpenStudy (anonymous):

now your job is to multiply \[\left(\sin(x)\cos(y)+\cos(x)\sin(y)\right)\times \sec(x)\sec(y)\]\]

OpenStudy (anonymous):

Maybe I should tell you this.. I've been gone from school for a few weeks now (death in the family) I have no idea how to do any of this..

OpenStudy (anonymous):

you get \[\sin(x)\cos(y)\sec(x)\sec(y)+\cos(x)\sin(y)\sec(x)\sec(y)\] as a first step

OpenStudy (anonymous):

okay.. then were the heck do you go?

OpenStudy (anonymous):

that was multiplying out using the distributive law then \(\sec(x)=\frac{1}{\cos(x)}\) so you can cancel some stuff

OpenStudy (anonymous):

lets look at the first term \[\sin(x)\cos(y)\sec(x)\sec(y)\] that is equal to \[\frac{\sin(x)\cos(y)}{\cos(x)\cos(y)}\]

OpenStudy (anonymous):

cancel the \(\cos(y)\) top and bottom and get \[\frac{\sin(x)}{\cos(x)}\] which you can rewrite as \[\tan(x)\]

OpenStudy (anonymous):

okay. that makes some sense..

OpenStudy (anonymous):

basically the same thing for the other side too then?

OpenStudy (anonymous):

did you know \(\sec(x)=\frac{1}{\cos(x)}\)? this is really algebra at this step

OpenStudy (anonymous):

yes, same thing for the second term

OpenStudy (anonymous):

only this time you will get \(\tan(y)\)

OpenStudy (anonymous):

Yes I did. Just didn't understand the problem.. so you get down to tan x=tanx?

OpenStudy (anonymous):

ohh, yeah. tan x + tan y

OpenStudy (anonymous):

that is it, yes

OpenStudy (anonymous):

So is that your answer then? tan x + tan y?

OpenStudy (anonymous):

that is my answer, yes

OpenStudy (anonymous):

Okay, that makes sense. Do you have time to help me with one more problem?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

okay, let me take a screen shot.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

I kind of understand it.. but not really :/

OpenStudy (anonymous):

ok we can do this, but it is going to take a minute ready?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

we are going to use the same formula that we used above \[\sin(u+v)=\sin(u)\cos(v)+\cos(u)\sin(v)\] at the moment, we know only two of those four numbers

OpenStudy (anonymous):

yes, we know sin u and cos v

OpenStudy (anonymous):

we need \(cos(u)\) do you know how to find it?

OpenStudy (anonymous):

Don't you have to make a triangle or something?

OpenStudy (anonymous):

yes exactly

OpenStudy (anonymous):

Okay, I'm not sure exactly how to come about finding it on a triangle

OpenStudy (anonymous):



OpenStudy (anonymous):

third side via pythagoras \[\sqrt{61^2-11^2}=60\]

OpenStudy (anonymous):

Wouldn't it be 50?

OpenStudy (anonymous):

no \[\cos(u)=\frac{60}{61}\]

OpenStudy (anonymous):

adjacent over hypotenuse

OpenStudy (anonymous):

Okay. so we still ned to find the sin one right?

OpenStudy (anonymous):

right but actually i made a mistake \[\cos(u)=-\frac{60}{61}\] because you are in quadrant 2

OpenStudy (anonymous):

now we need \(\sin(v)\) and it is the same idea as before 

OpenStudy (anonymous):

the third side is \[\sqrt{41^2-40^2}=9\]so \(\sin(v)=\frac{9}{41}\)

OpenStudy (anonymous):

now plug the numbers directly in to the formula

OpenStudy (anonymous):

okay, then you just solve it right? multiplying and adding fractions basically?

OpenStudy (anonymous):

\[\frac{11}{61}\times (-\frac{40}{41})+\frac{9}{61}\times(- \frac{60}{61})\]

OpenStudy (anonymous):

yeah it is arithmetic from here on in gotta run, good luck

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

wait quick

OpenStudy (anonymous):

oh i made a typo!!

OpenStudy (anonymous):

for multiplying you flip the second fraction?

OpenStudy (anonymous):

\[\frac{11}{61}\times (-\frac{40}{41})+\frac{9}{41}\times(- \frac{60}{61})\]

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

flip nothing multiply straight across

OpenStudy (anonymous):

okay. ill give it a try... thanks!

OpenStudy (anonymous):

yw